OpenStudy (anonymous):

Simplify: (9 – 8)2 × (24 – 6 /3)

OpenStudy (anonymous):

(1)2 X (18/3) = 2 X 6 = 12

OpenStudy (anonymous):

Always use PEMDAS!! (9 - 8)^2 (24 - 6 / 3) 9 - 8 = 1 1^2 = 1 24 - 6 /3 6/3 = 2 24 - 2 = 22 so 1 * 22 = 22 22 is your answer